OpenStudy (anonymous):

The revenue of a charter bus company depends on the number of unsold seats. If 100 seats are sold, the price is $55 per seat. Each unsold seat increases the price per seat by $1. Let X represents the number of seats that are unsold. a) Write an expression for the number of seats that are sold. b) Write an expression for the price per seat. c) Write an expression for the revenue d) Find the number of unsold seats that will produce the maximum revenue. e) Find the maximum revenue

OpenStudy (whpalmer4):

if \(x\) is the number of unsold seats, \(100-x\) is the number sold. Problem doesn't explicitly state that the bus seats 100, but it is implied. price per seat is $55 + $1*unsold seats = $55 +$1x revenue is price per seat * seats sold revenue should be a parabola. find the vertex of the parabola to find the number of seats unsold for maximum revenue find the revenue with that number of unsold seats

OpenStudy (anonymous):

can you please elaborate further.from c downwards.please

OpenStudy (whpalmer4):

No, that's the part you're supposed to do! I told you what the expression for seats sold is, and what the expression for the price per seat is, and that revenue is the product of those two expressions. Surely you can figure it out from that.

OpenStudy (anonymous):

so u mean the revenue is=$55 + $1 * 100-x?

OpenStudy (anonymous):

55 + 100- x?

OpenStudy (whpalmer4):

Revenue = (cost per seat) * (seats sold) cost per seat = $55 + $1x seats sold = 100-x What is (55+1x)(100-x)?

OpenStudy (anonymous):

ohh ok. so 55(100-x) + 1x(100-x) =5500-55x + 100x -x^2 5500+55x -x^2 im i on the right track?

OpenStudy (whpalmer4):

You need to recheck your algebra.

OpenStudy (anonymous):

ok right noow i have first sloved for the price per seat and i had 3080 + 56x

OpenStudy (whpalmer4):

No, you do not have a correct revenue equation. 5500 + 55x - x^2 is not correct. Fix that before you try to do something else.

OpenStudy (anonymous):

after trying to fix i got -56x^2 +2520x + 308000 for the revenue equation

OpenStudy (whpalmer4):

OFCOL. \[(55+1x)(100-x) = \]

OpenStudy (anonymous):

i got -x^2 + 55x + 5500

OpenStudy (whpalmer4):

No. write it out step by step. show me your work.

OpenStudy (whpalmer4):

If you keep giving the same answer, and someone keeps saying "do it again", that's a strong hint that maybe a different answer would be better :-)

OpenStudy (whpalmer4):

\[(55+1x)(100-x)\]apply distributive property \[55(100-x)+1x(100-x)\]apply distributive property again \[55*100 - 55*x + 1x *100 - 1x^2\]now simplify and collect like terms, carefully :-)

OpenStudy (anonymous):

ok.i first of all distributed the numbers 55(100-X) + X(100-X) 5500 - 55X +100X -X^2 5500 + 55X -X^2 =-X^2 + 55X + 5500

OpenStudy (whpalmer4):

On the other hand, if you feel that buying something for $55 should get you $55 change from $100, tell me where you are working as a cashier — I'd like to buy a whole bunch of stuff from you :-)

OpenStudy (whpalmer4):

-55X + 100X = ?

OpenStudy (anonymous):

OH MY GOODNESS ..HAHAHA

OpenStudy (whpalmer4):

:-)

OpenStudy (whpalmer4):

"watch out, forehead, here comes palm!" :-)

OpenStudy (anonymous):

OK SO -X^2 + 45X +5500

OpenStudy (anonymous):

LOL.THAT WAS AN OVERSIGHT

OpenStudy (whpalmer4):

yes, now we're talking! that's a parabola, right? do you know how to find the vertex?

OpenStudy (anonymous):

YES. LET ME WORK IT OUT AND TELL YPU MY ANSWE THEN,

OpenStudy (whpalmer4):

Sounds good.

OpenStudy (anonymous):

OK RIGHT NOW IM BEGINNING TO GET LOST AGAIN. WITH THE PARABOLA a=-1 b= 45 c= 550 right?

OpenStudy (anonymous):

where the vertex h =-b/2a and vertex k =c-ah^2 right?

OpenStudy (anonymous):

i meant c=5500

OpenStudy (whpalmer4):

c = 5500 so the vertex is at x = -b/2a = -45/-2 =

OpenStudy (anonymous):

yes i already got that to be 22.5.

OpenStudy (anonymous):

and i had k=4993.75

OpenStudy (whpalmer4):

Okay, how do we sell 22.5 tickets? :-)

OpenStudy (whpalmer4):

we have to go with 22 or 23, but which?

OpenStudy (whpalmer4):

(or not sell 22 or 23 tickets, I guess)

OpenStudy (anonymous):

lets go with 23

OpenStudy (anonymous):

so the 22.5 is the answer for d?

OpenStudy (whpalmer4):

well, we can't sell half a ticket, so we have to go with a number that is an integral number of tickets, right? as it turns out, the result is symmetrical — revenue is the same for 22 or 23 unsold tickets. it's just that the vertex of the parabola is slightly higher, if we could only sell half tickets :-)

OpenStudy (whpalmer4):

now who knows, maybe the person who wrote the question expects the nonsensical answer of 22.5...I would make a note explaining that 22.5 is the peak of the curve, but does not represent a real-life solution. what is the revenue at 22 or 23 unsold tickets?

OpenStudy (anonymous):

oh ok .lets say i go with 23 ryt? too find e) do i would need to substitute the 23 in the revevue equation i had.thus the -X^2 + 45X +5500 to get the maximum revenue?

OpenStudy (whpalmer4):

yes, exactly. -(23)^2 + 45(23)+5500 and try it with 22 as well, just to convince yourself that it is true, although given that we know parabolas are symmetric about the vertex and our vertex is exactly midway between 22 and 23, it shouldn't be too surprising.

OpenStudy (whpalmer4):

it looks like we only lose $0.25 for not being able to sell fractional tickets — hardly worth the trouble :-)

OpenStudy (anonymous):

oh ok .thanks very much one more question.

OpenStudy (anonymous):

Application of Quadratic Equations Business The manager of a bicycle shop has found that, at a price (in dollars) of p(x)=150-x/5 per bicycle, x bicycles will be sold. a)Find an expression for the total revenue from the sale of x bicycles. (Hint: Revenue = Demand X Price.) b)Find the number of bicycle sales that leads to maximum revenue. c)Find the maximum revenue.

OpenStudy (whpalmer4):

what did you get for revenue?

OpenStudy (anonymous):

i havent worked for it since i know how to go about it..:-)

OpenStudy (whpalmer4):

no, for the last question!

OpenStudy (anonymous):

using 23 i had 6006

OpenStudy (anonymous):

i had the same for 22

OpenStudy (whpalmer4):

Okay, good.

OpenStudy (anonymous):

ok now please help me with the second question